OpenStudy (anonymous):

solve the equations: 1. 3^(x+1) =7 2. 5^(x+1) = 25^(1-x) we are learning logarithm right now..

OpenStudy (anonymous):

\[3^{x+1}=7\] apply log on both sides \[log 3^{x+1}=log 7\] \[(x+1) log 3=log 7\] \[(x+1) =log 7/ log 3\]

OpenStudy (anonymous):

x = (log7/log3)-1

OpenStudy (anonymous):

do the other problem in a similar manner

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you are welcome